1) find the distance between the following

a) (4,7) and (3,7)

2) write the equation of line with slope (-2/3)and passing through the points ( -2,5)


1
Expert's answer
2021-07-12T16:22:52-0400

1) "d=\sqrt{(4-3)^2+(7-7)^2}=1."


2) "y-5=-\frac{2}{3}(x+2)" or "y=-\frac{2}{3}x+\frac{11}{3}."